# Junsheng Huang1,2, Zhitao He1, Sandeep Polisetty3, Qingyun Wang3, May Fung1,2

### v1

@@ Abstract
With the widespread application of large language models (LLMs), the issue of generating non-existing facts, known as hallucination, has garnered increasing attention. Previous research in enhancing LLM confidence estimation mainly focuses on the single problem setting. However, LLM awareness of its internal parameterized knowledge boundary under the more challenging multi-problem setting, which requires answering multiple problems accurately simultaneously, remains underexplored. To bridge this gap, we introduce a novel method, Multiple Answers and Confidence Stepwise Tuning (MAC-Tuning), that separates the learning of answer prediction and confidence estimation during fine-tuning on instruction data. Extensive eperiments demonstrate that our method proposed outperforms baselines by up to 25% in average precision. [^1] Our code and resource will be released upon publication.
@@
Junsheng Huang1,2, Zhitao He1, Sandeep Polisetty3, Qingyun Wang3, May Fung1,2 1Hong Kong University of Science and Technology 2University of Illinois     3UMass Amherst jh103@illinois.edu      yrfung@ust.hk
@@ 1 Introduction
Large language models (LLMs) are widely used in knowledge-intensive scenarios, such as question answering (Gu et al., 2023), information retrieval (Ren et al., 2023), and recommendation systems (Liu et al., 2023). Yet, they often produce non-existing facts when faced with questions outside their parametric knowledge, which undermines their reliability (Maynez et al., 2020). Many efforts have been dedicated to mitigating LLM hallucination, such as leveraging knowledge boundaries to constrain the reasoning scope of LLMs to help them better distinguish between reliable and unreliable information Chen et al. (2024); Liang et al. (2024a); Zhang et al. (2024). Notably, these work mainly focus on the single-problem setting, where users repeatedly input questions and context for models to answer one by one.
@@
LLM hallucination in the multi-problem setting — in which a single input contains multiple distinct sub-questions for the model to extract and address — remains relatively underexplored. As seen in Figure 1, this is a fundamentally challenging setting because the language model must distinguish each sub-question, reason over different knowledge, and synthesize results cohesively. Undesirable overshadowing of context from one sub-question with another, and propagation of reasoning confusion, may compromise the reliability of LLMs in multi-problem answering (Cheng et al., 2023, Wang et al., 2024, Son et al., 2024). As LLM-based multi-problem reasoning becomes increasingly widespread due to its efficiency benefits in scenarios involving extensive shared contexts (e.g., task instructions, exemplars), reduced model access, and lower API costs, enhancing model confidence estimation calibration for this emerging class of reasoning demands growing attention and effort as well.
In this paper, we investigate the hallucinations in LLMs within the multi-problem setting and propose leveraging the knowledge boundary to simultaneously handle the composition of multiple problems. Inspired by Zhang et al. (2024), which advocates for encouraging the LLM to express confidence to reduce hallucinations, we introduce Multiple Answers and Confidence Stepwise Tuning (MAC-Tuning) under multi-problem setting. Our approach involves several key steps. First, we identify the knowledge boundary between parametric knowledge and the multi-problem dataset to extract uncertain questions. Next, we automatically label the model’s confidence for both certain and uncertain data. These labeled data are then used to create multiple question-answer data and multiple QA-Confidence data so we can train the original model by separating the learning process of ground-truth answers and confidence, which enhances performance and reliability.
@@
  • Through extensive experiments across various settings and datasets, MAC-Tuning achieves an AP score gain of up to 25% over baselines in LLM multi-problem reasoning. Finally, we share our insights discovered to motivate future work.
@@ 2.1 Multi-Problem Tuning Data Construction
First, we randomly combine n single problems from original datasets to construct our initial Multi-Problem dataset. We utilize this to compare LLMs’ outputs with ground-truth answers, for distinguishing the knowledge boundary between LLM parameters and instruction data. Specifically, for each individual problem in the multi-problem pair, we assign: “I am sure” if the model output aligns with ground-truth answer; “I am unsure” elsewise (e.g., Step 2 in Figure 2). With the assigned confidence labels, we construct Multi-Problem Tuning data as follows:
Multiple QA pair $D_{MultQA}$: We directly combine the questions and answers together, with Question $q_{i}$ as input and Answer $a_{i}$ as output label, to form $D_{MultQA}=[(q_{1},a_{1})...(q_{i},a_{i})...(q_{n},a_{n})]$.
@@ 2.2 Training and Inference
where $Q$, $A$, and $C$ represent the sets of multiple questions, multiple answers, and multiple confidence levels, respectively. $\Theta_{0}$ and $\Theta_{1}$ represent the parameters of the base model and the model after the first step of fine-tuning, respectively.
@@ 3.6 Analysis on Various Number of Questions
We explore different numbers of questions in the multi-problem setting to investigate how this varies the accuracy. We only do this for three Independent setting datasets, and the results are reported in Figure 3. MAC-Tuning consistently outperforms the base model in accuracy by at least 10.0% and, on average, 26.8%. For easy tasks like ParaRel, The ability of the base model to handle multiple problems simultaneously is even higher when compared with the traditional single-problem setting, indicating that LLM could leverage in-context learning and focus on relevant knowledge better under multi-problem setting. However, for other datasets like MMLU, MAC-Tuning performs slightly worse as the question number increases. A reasonable explanation is that it is out of the base model’s ability to learn too many hard tasks together but within effective scope to learn several easy tasks at the same time. Further studies are detailed in A.9.
Figure 3: Accuracy for combining different number (n) of single problem together. Solid lines represent MAC Tuning, while dashed lines represent LLaMA3.
3.7 Analysis on Different Base Model Choice
@@
Table 4: Confidence calibration result (%) for Qwen2-7B-Instruct, with bold denoting the top performance across different methods.
@@ A.2 Related Work
Hallucination: Large language models (LLMs) are widely used in knowledge-intensive scenarios, such as question answering (Gu et al., 2023), information retrieval (Ren et al., 2023) and recommendation systems (Liu et al., 2023). However, LLMs have tendency to generate non-existing facts when faced with questions that are out of their parametric knowledge (Maynez et al., 2020). Many efforts are dedicated to mitigating hallucinations in LLMs, such as retrieval-augmented generation (Gao et al., 2024,Peng et al., 2023), multi-agent debate (Du et al., 2023,Sun et al., 2023), and model confidence calibration (Zhang et al., 2024, Hu et al., 2023).
@@
Multiple Problem Setting: Current LLM research has predominantly focused on single problem setting. There are only a few works focusing on this new setting. (Cheng et al., 2023) propose batch prompting that prompts LLMs with single independent problems batched together following few-shot exemplars together. Son et al. (2024) goes further by researching sequential datasets and develops the first multi-task benchmark (MTI-Bench). Wang et al. (2024) pays attention to zero-shot cases of multi-problem setting and design a new benchmark ZeMPEB. Overall, LLM hallucination under this multi-problem setting is still under-explored.
@@ A.4 Dataset Details
  • MMLU (Hendrycks et al., 2021) a dataset covering different subjects and difficulty. It tests both world knowledge and problem solving ability, which has good granularity and breath. We directly use the modified dataset from Zhang et al. (2024) in our Multiple Choice setting.
  • CoQA (Reddy et al., 2019) a dataset designed to evaluate the ability of models to understand and generate answers in a conversational setting. We randomly pick 5k training dataset from theirs. In Question Answer setting, we combine multiple questions together under the same "story" category in the dataset.
  • MTI Bench (Son et al., 2024) a comprehensive evaluation benchmark encompassing 5,000 instances across 25 tasks. We pick the sequential part of this benchmark and divide it into 800 training data and 200 test data.
  • SQA (Iyyer et al., 2017) a dataset designed to explore the task of answering sequences of inter-related questions on HTML tables. We pick 5 sequential questions for each HTML table and have 3985 training data.
@@ A.8 Detailed Information for Variant Methods
A.9 Cross Task Transfer Study
We fine-tune the model with question number n = 3 and let it response to question number n = 1 (which is Single Problem setting) and question number n = 5. For the former one, we want to test if the model understands the single problem. For the latter one, we want to test if MAC-Tuning can generalize across different question number n. The results are reported in Table 5.
Question Number CoQA ParaRel GSM MMLU
n = 1 78.78 84.24 71.12 54.60
n = 5 79.12 86.16 67.67 63.69
Table 5: Accuracy (%) for MAC-Tuning with question number n = 3 transferring to question number n = 1 (which is single-problem setting) and question number n = 5. We use one-shot CoT for GSM results.
From the result of n = 1, we observe that accuracy improves in easy dataset like CoQA but decreases in difficult dataset like GSM, comparing with using question number n = 1 to fine-tune. This indicates that LLM learn knowledge during fine-tuning rather than just memorizing the patterns for multi-problem setting. With n = 5, the model performs about the same or even better than using question number n = 5 to fine-tune. This strengthen the statement we make in Section 3.6 that LLMs are more efficient at learning multiple easy tasks but struggle with several difficult tasks simultaneously.

### v4

# Junsheng Huang1,2, Zhitao He1, Yuchen Huang1, Sandeep Polisetty3, Qingyun Wang4 Yi R. (May) Fung1

@@ Abstract
The hallucination of non-existent facts by LLMs is an important problem given its widespread adoption across various applications. Previous research addresses this problem by analyzing the internal parameterized knowledge boundaries to estimate confidence. However, these studies focus on the single-problem setting and have not explored the more challenging multi-problem setting, which requires accurately answering multiple questions simultaneously. We introduce a novel method for the multi-problem setting, Multiple Answers and Confidence Stepwise Tuning (MAC-Tuning), that separates the learning of answer prediction and confidence estimation during fine-tuning on instruction data. Extensive experiments demonstrate that our method outperforms baselines by up to 25% in average precision. [^1] We release our code and resource at MAC-Tuning.
@@
Junsheng Huang1,2, Zhitao He1, Yuchen Huang1, Sandeep Polisetty3, Qingyun Wang4 Yi R. (May) Fung1 1Hong Kong University of Science and Technology 2University of Illinois Urbana-Champaign   3UMass Amherst  4William & Mary jh103@illinois.edu      yrfung@ust.hk
@@ 1 Introduction
Large language models (LLMs) are widely used in knowledge-intensive scenarios, such as question answering (Gu et al., 2023), information retrieval (Ren et al., 2023), and recommendation systems (Liu et al., 2023). Yet, they often produce non-existing facts when faced with questions outside their parametric knowledge, which undermines their reliability (Maynez et al., 2020; He et al., 2025c). Many efforts have been dedicated to mitigating LLM hallucination, such as leveraging knowledge boundaries to constrain the reasoning scope of LLMs to help them better distinguish between reliable and unreliable information Chen et al. (2024); Liang et al. (2024a); Zhang et al. (2024); Jin et al. (2024). Notably, these work mainly focus on the single-problem setting, where users repeatedly input questions and context for models to answer one by one.
@@
LLM hallucination in the multi-problem setting — in which a single input contains multiple distinct sub-questions with optional context for the model to extract and address — remains relatively underexplored. As seen in Figure 1, this is a fundamentally challenging setting because the model must distinguish each sub-question, reason over different knowledge, and synthesize results cohesively. Undesirable overshadowing of context from one sub-question with another, and propagation of reasoning confusion, may compromise the reliability of LLMs in multi-problem answering (Cheng et al., 2023a, Wang et al., 2024, Son et al., 2024; Li et al., 2024; He et al., 2025b). As LLM-based multi-problem reasoning becomes increasingly widespread due to its efficiency benefits in scenarios involving extensive shared contexts (e.g., task instructions, exemplars), reduced model access, and lower API costs, enhancing model confidence estimation calibration for this emerging class of reasoning demands growing attention and effort as well.
In this paper, we investigate the hallucinations in LLMs within the multi-problem setting and propose leveraging the knowledge boundary to simultaneously handle the composition of multiple problems. Inspired by Zhang et al. (2024), which advocates for encouraging the LLM to express confidence to reduce hallucinations, we introduce Multiple Answers and Confidence Stepwise Tuning (MAC-Tuning) under the multi-problem setting. Our approach involves several key steps. First, we identify the knowledge boundary between parametric knowledge and the multi-problem dataset to extract uncertain questions. Next, we automatically label the model’s confidence for both certain and uncertain data. These labeled data are then used to create multiple question-answer data and multiple QA-Confidence data so we can train the original model by separating the learning process of ground-truth answers and confidence, which enhances performance and reliability.
@@
  • Through extensive experiments with different base models of varying sizes and various datasets, MAC-Tuning achieves an AP score gain of up to 25% over baselines in LLM multi-problem reasoning. Finally, we share our insights discovered to motivate future work.
@@ 2.1 Multi-Problem Tuning Data Construction
First, we combine n single problems from original datasets to construct our initial Multi-Problem dataset. We utilize this to compare LLMs’ outputs with ground-truth answers, for distinguishing the knowledge boundary between LLM parameters and instruction data. Specifically, for each individual problem in the multi-problem pair, we assign: “I am sure” if the output aligns with ground-truth answer; “I am unsure” elsewise (e.g., Step 2 in Figure 2). With the assigned confidence labels, we construct Multi-Problem Tuning data as follows: Multiple QA pair $D_{MultQA}$: We directly combine the questions and answers together, with Question $q_{i}$ as input and Answer $a_{i}$ as output label, to form $D_{MultQA}=[(q_{1},a_{1})...(q_{i},a_{i})...(q_{n},a_{n})]$.
@@ 2.2 Training and Inference
where $Q$, AA, and $C$ represent the sets of multiple questions, multiple answers, and multiple confidence levels, respectively. $\Theta_{0}$ and $\Theta_{1}$ represent the parameters of the base model and the model after the first step of fine-tuning, respectively.
@@ 3.6 Analysis on Various Number of Questions
We explore different numbers of questions in the multi-problem setting to investigate how this varies the accuracy. We only do this for three Independent setting datasets, and the results are reported in Figure 3. MAC-Tuning consistently outperforms the base model in accuracy by at least 10.0% and, on average, 26.8%. For easy tasks like ParaRel, the ability of the base model to handle multiple problems simultaneously is even higher when compared with the traditional single-problem setting, indicating that LLM could leverage in-context learning and focus on relevant knowledge better under the multi-problem setting. However, for other datasets like MMLU, MAC-Tuning performs slightly worse as the question number increases. A reasonable explanation is that it is out of the base model’s ability to learn too many hard tasks together but within effective scope to learn several easy tasks at the same time. The result for extremely large numbers of questions is in Appendix A.9.
Figure 3: Accuracy for combining different number (n) of single problem together. Solid lines represent MAC-Tuning, while dashed lines represent LLaMA3.
For future work, we believe it is meaningful to further explore whether mixing questions of varying difficulty and diversity in multi-problem settings leads to better scaling behavior Qin et al. (2025). This direction may help uncover strategies for enhanced model generalization.
3.7 Cross Task Transfer Study
We fine-tune the model with question number n = 3, and subsequently evaluate its performance on both single problem inputs (n = 1) and more complex instances involving a higher number of questions (n = 5). The evaluation on n = 1 aims to examine whether the model retains its ability to accurately solve individual problems after being exposed to multi-problem setting training. Conversely, the evaluation on n = 5 serves to assess the generation capability of MAC-Tuning when scaling to larger compositions beyond the training scope. This allows us to understand both the robustness and scalability of our proposed method across different levels of compositional complexity. The results are reported in Table 4.
Question Number CoQA ParaRel GSM MMLU
n = 1 78.8 84.2 71.1 54.6
n = 5 79.1 86.2 67.7 63.7
Table 4: Accuracy (%) for MAC-Tuning with question number n = 3 transferring to question number n = 1 and question number n = 5. We use one-shot CoT for GSM results.
From the result of single question inputs, we observe that accuracy increases on easier dataset (e.g. CoQA) but decreases on more challenging dataset like GSM, when compared to models fine-tuned specifically under that setting. This pattern indicates that LLM acquires underlying knowledge during MAC-Tuning rather than merely memorizing the patterns for multi-problem setting. In contrast, when fine-tuned on five-question inputs (n = 5), the model’s performance is comparable or even exceeds that of the baseline fine-tuned directly on n = 5. These findings strengthen the statement we make in Section 3.6: while LLMs can efficiently learn multiple easy tasks, they exhibit difficulty when faced with several difficult tasks simultaneously.
3.8 Analysis on Different Base Model
@@
Table 5: Confidence calibration result (%) for Qwen2-7B-Instruct, with bold denoting the top performance.
3.9 Analysis on Different Model Size
We compare base models of different sizes to study how model size affects performance and the confidence calibration results for Llama-3.2-3B (Dubey et al., 2024) and Phi-3.5-mini-Instruct (Abdin et al., 2024) is shown in Table 6 and Table 7 respectively. Despite using different base models of varying sizes, the results indicate consistent performance patterns. For smaller models, the accuracy improvement after MAC-Tuning is more evident, indicating enhanced ability to differentiate between certain and uncertain questions.
 Independent  Sequential
Approach   ParaRel    CoQA MTI-Bench        SQA
AP ECE AP ECE AP ECE AP ECE
Vanilla 30.7 70.3 46.0 45.0 34.3 70.3 35.6 45.0
MAC-Tuning 55.5 33.5 62.4 33.5 35.5 33.5 44.3 33.5
Table 6: Confidence calibration result (%) for Llama-3.2-3B, with bold denoting the top performance across different methods.
 Independent  Sequential
Approach   ParaRel    CoQA MTI-Bench        SQA
AP ECE AP ECE AP ECE AP ECE
Vanilla 58.0 22.8 56.0 32.9 21.4 29.1 96.6 33.7
MAC-Tuning 70.2 14.2 68.2 29.0 68.7 22.6 52.3 23.9
Table 7: Confidence calibration result (%) for Phi-3.5-mini-Instruct, with bold denoting the top performance across different methods.
3.10 Human Evaluation
We randomly selected and evaluated 100 examples from the ParaRel dataset. The human annotator was shown only the query and the ground-truth answer, and asked to assess the factual correctness of each model’s output without knowing the confidence label (“I am sure” or “I am unsure”). We then compared the accuracy rates between responses that the model labeled as “I am sure” versus those labeled as “I am unsure.” For answers the model labeled as “I am sure”: Human evaluation confirmed a factual accuracy of 89.2%. For answers the model labeled as “I am unsure”: The human-verified factual accuracy was only 41.2%.
These results provide strong empirical evidence for our central claim. The substantial 48-percentage-point gap demonstrates that the confidence learned by MAC-Tuning is not merely an artifact of automatic metrics. Instead, it reflects a genuine, human-perceptible distinction in answer quality. This strong alignment with human judgment validates the reliability and real-world applicability of our method.
@@ A.2 Related Work
Hallucination: Large language models (LLMs) are widely used in knowledge-intensive scenarios, such as question answering (Gu et al., 2023; He et al., 2025c), information retrieval (Ren et al., 2023; Huang et al., 2025) and recommendation systems (Liu et al., 2023). However, LLMs have tendency to generate non-existing facts when faced with questions that are out of their parametric knowledge (Maynez et al., 2020). Many efforts are dedicated to mitigating hallucinations in LLMs, such as retrieval-augmented generation (Gao et al., 2024,Peng et al., 2023), multi-agent debate (He et al., 2023; Du et al., 2023; He et al., 2024; Sun et al., 2023; He et al., 2025a), and model confidence calibration (Zhang et al., 2024; Hu et al., 2023; He et al., 2025c).
@@
Multiple Problem Setting: Current LLM research has predominantly focused on single problem setting. There are only a few works focusing on this new setting. Cheng et al. (2023a) propose batch prompting that prompts LLMs with single independent problems batched together following few-shot exemplars together. Son et al. (2024) goes further by researching sequential datasets and develops the first multi-task benchmark (MTI-Bench). Wang et al. (2024) pays attention to zero-shot cases of multi-problem setting and design a new benchmark ZeMPEB. Li et al. (2024) analyze different strategy under independent setting, where single questions are combined into various constraint formats without sharing context between them. Despite these efforts, the multi-problem setting presents significant challenges. For example, Wang et al. (2024) shows that in zero-shot setting, LLMs consistently perform worse when selecting indices of texts for a given class label with multiple mixed-source reasoning problems. Similarly, for few-shot setting, Cheng et al. (2023b) and Lin et al. (2024) have found that the overall accuracy decreases with the increase in batch size. Notably, this setting is also meaningful in real-world applications: for independent scenario, batching unrelated queries can reduce model calls and API costs; for sequential scenario, where questions share context—such as in math problem solving, data processing, or software debugging—the correctness of each intermediate reasoning step is critical. Overall, hallucination and performance instability under the multi-problem setting are still under-explored and present significant challenges for current LLMs.
@@ A.4 Dataset Details
  • MMLU (Hendrycks et al., 2021): a dataset covering different subjects and difficulty. It tests both world knowledge and problem solving ability, which has good granularity and breath. We directly use the modified dataset from Zhang et al. (2024) in our Multiple Choice setting.
  • CoQA (Reddy et al., 2019): a dataset designed to evaluate the ability of models to understand and generate answers in a conversational setting. We randomly pick 5k training dataset from theirs. In Question Answer setting, we combine multiple questions together under the same "story" category in the dataset.
  • MTI Bench (Son et al., 2024): a comprehensive evaluation benchmark encompassing 5,000 instances across 25 tasks. We pick the sequential part of this benchmark and divide it into 800 training data and 200 test data.
  • SQA (Iyyer et al., 2017): a dataset designed to explore the task of answering sequences of inter-related questions on HTML tables. We pick 5 sequential questions for each HTML table and have 3985 training data.
@@ A.8 Detailed Information for Variant Methods
A.9 Extremely Large Multi-problem Scenarios
We also conduct experiments to include larger values of question numbers (n = 10 and 15). These results are shown in table 8.
Pararel n=1 n=2 n=3 n=4 n=5 n=10 n=15
LLaMA3 42.9 45.3 42.2 47.0 48.2 48.2 49.6
MAC-Tuning 86.1 81.6 88.0 83.7 86.3 82.6 77.1
Table 8: Accuracy (%) comparison between LLaMA3 and MAC-Tuning on ParaRel dataset with extremely large question number of combined questions $n$.
The extended results show that MAC-Tuning consistently and significantly outperforms the Llama3 baseline across all tested values of n. Notably, even at n = 15, our method achieves 77.1% accuracy, maintaining a substantial margin of over 27.5 percentage points against the baseline (49.6%).
One notable point is that for different values of n, we still use 3 training epochs for fairness. However, we only have around 5000 training questions for the Pararel dataset. Thus, for a large value of n (taking n = 10 as an example), each epoch only has 500 training data and the model might be under training because of this.
Overall, it provides a clearer picture of our method’s performance scalability and demonstrates its robustness in handling more complex, large-scale multi-problem inputs.